### v1

# Spatial Conformal Inference through Localized Quantile Regression

Abstract
Reliable uncertainty quantification at unobserved spatial locations, especially in the presence of complex and heterogeneous datasets, remains a core challenge in spatial statistics. Traditional approaches like Kriging rely heavily on assumptions such as normality, which often break down in large-scale, diverse datasets, leading to unreliable prediction intervals. While machine learning methods have emerged as powerful alternatives, they primarily focus on point predictions and provide limited mechanisms for uncertainty quantification. Conformal prediction, a distribution-free framework, offers valid prediction intervals without relying on parametric assumptions. However, existing conformal prediction methods are either not tailored for spatial settings, or existing ones for spatial data have relied on rather restrictive i.i.d. assumptions. In this paper, we propose Localized Spatial Conformal Prediction (LSCP), a conformal prediction method designed specifically for spatial data. LSCP leverages localized quantile regression to construct prediction intervals. Instead of i.i.d. assumptions, our theoretical analysis builds on weaker conditions of stationarity and spatial mixing, which is natural for spatial data, providing finite-sample bounds on the conditional coverage gap and establishing asymptotic guarantees for conditional coverage. We present experiments on both synthetic and real-world datasets to demonstrate that LSCP achieves accurate coverage with significantly tighter and more consistent prediction intervals across the spatial domain compared to existing methods.
Keywords: conformal prediction, uncertainty quantification, spatial statistics, mixing
1 Introduction
Quantifying uncertainty at unobserved spatial locations has been a longstanding challenge in spatial statistics, particularly in practical applications such as weather forecasting (Siddique et al., 2022) and mobile signal coverage estimation (Jiang et al., 2024). Traditional methods like Kriging rely on strong parametric assumptions, including normality and stationarity, to model spatial relationships and quantify uncertainty (Cressie, 2015). However, the failure of these assumptions in the complex spatial datasets (Heaton et al., 2019) results in unreliable uncertainty quantification. The issue is especially pronounced when constructing prediction intervals, as deviations from stationarity or Gaussianity can critically undermine their validity (Fuglstad et al., 2015).
While many methods have been developed to handle heterogeneity (Gelfand et al., 2005; Duan et al., 2007), these approaches are often computationally expensive and cannot scale effectively for massive datasets. Furthermore, fully modeling the underlying process is not always necessary, particularly when the goal is to produce reliable prediction intervals. Recently, machine learning approaches have offered alternative strategies for spatial prediction (Hengl et al., 2018; Chen et al., 2020), though they tend to focus on point predictions and often lack rigorous uncertainty quantification.
Conformal prediction, introduced by (Vovk et al., 2005), provides a powerful, distribution-free approach to uncertainty quantification. Its ability to generate valid prediction sets without assumption on the underlying data distribution and the prediction model has gained widespread popularity in both machine learning and statistics (Lei & Wasserman, 2014; Angelopoulos et al., 2023). By leveraging only the exchangeability of data, conformal prediction ensures valid coverage at any significance level, making it highly attractive for scenarios where a black-box model is used, or traditional parametric assumptions may fail.
However, in many real-world datasets—such as time-series data—the assumption of exchangeability does not hold. To address this, recent work has extended conformal prediction to handle non-exchangeable data. For instance, (Tibshirani et al., 2019) introduced weighted quantiles to maintain valid coverage in the presence of distributional shifts between training and test sets by leveraging the likelihood ratio between distributions. More recently, (Barber et al., 2023) tackled the challenge of distribution shifts by bounding the coverage gap using the total variation distance, although the issue of optimizing these weights remains open. For time-series data, further improvements have been made in tightening prediction intervals and building theoretical guarantees, as demonstrated by (Xu & Xie, 2023; Xu et al., 2024).
In this paper, we extend conformal prediction to spatial data, where the assumption of exchangeability rarely holds. While time-series data can be viewed as a special case of spatial data defined in a one-dimensional time domain, spatial data is inherently multidimensional and poses unique challenges. For example, while time indices are typically discrete and naturally ordered, spatial locations are continuous and lack intrinsic ordering. Despite the prevalence of spatial data in real-world applications, there has been limited work on conformal prediction methods tailored to this context. To address this gap, we propose Localized Spatial Conformal Prediction (LSCP), a novel conformal prediction method that employs localized quantile regression for constructing prediction intervals. Our method and theoretical framework can also be extended to spatio-temporal settings, broadening its applicability. Here is a revised version of the summary:
  • Localized Spatial Conformal Prediction (LSCP): We introduce LSCP, a conformal prediction algorithm specifically designed for spatial data, which utilizes localized quantile regression to construct prediction intervals.
  • Theoretical guarantees: We establish a finite-sample bound for the coverage gap and provide asymptotic convergence guarantees for LSCP, without requiring the exchangeability of the data.
  • Numerical Evaluation: We extensively evaluate LSCP against state-of-the-art conformal prediction methods using both synthetic and real-world datasets. The results highlight LSCP’s ability to achieve tighter prediction intervals with valid coverage and more consistent performance across the spatial domain.
1.1 Literature
Conformal prediction beyond exchangeability.
Traditional conformal prediction relies on the assumption of i.i.d. or exchangeability, which is often difficult to satisfy in real-world datasets. Recent research has focused on extending methods and theoretical frameworks to non-exchangeable data to broaden applicability. A prominent approach is weighted conformal prediction, which assigns greater importance to samples deemed more ”reliable.” The work from (Tibshirani et al., 2019) explored the covariate shift scenario, where the distribution of features $X$ differs between calibration and test data, while the conditional distribution $Y|X$ remains unchanged. They demonstrated that weighting samples by the ratio of the distributions restores exchangeability. Building on this, (Barber et al., 2023) proposed a general weighted conformal framework and provided an analysis of the coverage gap in the general non-exchangeable setting, offering insights into the relationship between the weights and the coverage gap.
Conformal prediction for time series.
One key application of conformal prediction in non-exchangeable settings is time-series data. Some studies (Gibbs & Candes, 2021; Zaffran et al., 2022) have focused on adaptively adjusting the significance level $\alpha$ over time to achieve valid coverage. Another work (Angelopoulos et al., 2024), built upon the idea of control theory, prospectively model the non-conformity scores in an online setting. Another line of research (Tibshirani et al., 2019; Xu et al., 2024) follows the concept of weighted conformal prediction by assigning higher weights to more recent data points. The choice of the weights plays a critical role in determining the empirical performance of the method. However, as noted by (Barber et al., 2023), no universally optimal weighting strategy has been found, leaving room for further exploration and optimization in this area. Without exchangeability in the time series setting, the finite-sample guarantee does not necessarily hold, and asymptotic coverage can be achieved instead with certain additional assumptions.
Conformal prediction for spatial data.
The spatial setting represents a broader and more complex domain compared to time-series data, yet research on conformal prediction for spatial contexts remains limited. A recent study by (Mao et al., 2024) introduced a spatial conformal prediction method under the infill sampling framework, where data density increases within a bounded region. The key finding is that for spatial data, the exact or asymptotic exchangeability holds in certain settings. The algorithm falls within the category of weighted conformal prediction, employing kernel functions such as the radial basis function (RBF) kernel as weights. Similarly, (Guan, 2023) proposed a general localized framework for conformal prediction. While not explicitly designed for spatial data, their method also advocates using kernel functions for weighting, highlighting its potential relevance in spatial applications.
2 Problem setup
In this paper, we consider a spatial setting with observations $\{Z(s_{i})\}_{i=1}^{n}$, where $Z(s)=(X(s_{i}),Y(s_{i}))$ represents a random field observed at a finite set of spatial locations $s_{i}$. Here, $Y(s)\in\mathbb{R}$ denotes the response variable, and $X(s)\in\mathbb{R}^{p}$ is the associated feature vector. The feature vector $X(s)$ can include any relevant information that is useful for predicting $Y(s)$, such as spatial location $s$. Unlike the time series setting, where observations are indexed by fixed times $t$, the spatial setting considers the locations $s$ as random variables sampled from a distribution $g(s)$, while allowing for potential dependency within the random field $Z(s)$.
In the context of conformal prediction, the objective is to construct a prediction region $\hat{C}_{n}(X(s_{n+1}))$ for an unobserved response $Y(s_{n+1})$ given a known feature vector $X(s_{n+1})$. For a user-specified confidence level $\alpha$, we aim to ensure that the probability of $Y(s_{n+1})$ falling within the prediction region exceeds $1-\alpha$. This notion of coverage can be interpreted in two ways: marginal coverage and conditional coverage. Marginal coverage is defined as
$\mathbb{P}(Y(s_{n+1})\in\hat{C}_{n}(X(s_{n+1})))\geq 1-\alpha,$
whereas conditional coverage requires that
$\mathbb{P}(Y(s_{n+1})\in\hat{C}_{n}(X(s_{n+1}))\mid X(s_{n+1}))\geq 1-\alpha.$
Conditional coverage is a stronger condition than marginal coverage as it requires valid coverage for different $X(s)$. However, as shown by (Foygel Barber et al., 2021), achieving conditional coverage universally is impossible without making additional assumptions about the data distribution. In traditional conformal prediction settings, where data points are assumed to be i.i.d. or exchangeable, only marginal coverage is typically guaranteed. Besides valid coverage, constructing a prediction region that is as narrow as possible is desirable to improve the empirical performance of conformal prediction. The whole space is an example of a trivial prediction set, but it is too large to be useful and informative.
3 Method
In this section, we will first introduce the background of conformal prediction and our proposed Local Spatial Conformal Prediction (LSCP) method. Then we will discuss the benefits of LSCP compared to several alternatives.
3.1 Background: conformal prediction
conformal prediction is a widely used technique for constructing prediction intervals with finite-sample guarantees under minimal assumptions. The method is designed to provide a prediction interval for a response variable $Y_{n+1}$ associated with a new feature vector $X_{n+1}$, given a predictive model $\hat{f}$ and a set of observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$. The data is divided into two disjoint sets: a training set used to fit the model $\hat{f}$, and a calibration set used to calculate non-conformity scores, which quantify the uncertainty of the model at each point. A common choice for non-conformity score is $\hat{\varepsilon}_{i}=|Y_{i}-\hat{f}(X_{i})|$, which measures how well the model prediction aligns with the observed response. There is no restriction on the choice of the non-conformity score.
To construct a prediction interval for $Y_{n+1}$, the empirical quantile of the non-conformity scores $\hat{Q}_{n}$ from the calibration set is used as the estimate. Specifically, for a user-specified confidence level $1-\alpha$, the prediction interval is defined as
$\hat{C}_{n}(X_{n+1})=\left[\hat{f}(X_{n+1})-\hat{Q}_{n}(1-\alpha),\;\hat{f}(X_{n+1})+\hat{Q}_{n}(1-\alpha)\right].$
The empirical quantile can be written explicitly as
$\hat{Q}_{n}(p)=\inf\{e\in\mathbb{R}:\sum_{i=1}^{n}\frac{1}{n}\mathbf{1}\{\hat{\varepsilon}(s_{i})\leq e\}\leq p\}.$
Utilizing the exchangeability of the data, the method provides a prediction interval with valid marginal coverage, ensuring that $Y_{n+1}$ falls within the interval with probability at least $1-\alpha$. Besides the strong theoretical guarantee, conformal prediction is appealing due to its flexibility in the prediction model, as it has no assumptions about the underlying distribution of $Y$ or the form of the model $\hat{f}$.
Classical geostatistical methods like Gaussian Process (GP) assume that any finite set of observations has a joint Gaussian distribution. Given data, it provides both mean prediction and uncertainty estimates for unseen points, making it suitable for regression tasks where probabilistic predictions are desirable. In contrast to conformal prediction, which constructs prediction intervals by assessing model residuals without assumptions on data distribution, GPs rely on a Gaussian prior and explicit covariance structure. While conformal prediction provides finite-sample coverage guarantees under minimal assumptions, GPs often lose coverage when data does not satisfy the Gaussian assumption. Furthermore, it is computationally expensive to learn a GP model, which makes it difficult for GP to utilize as much data as possible. This also weakens the performance of GP. On the contrary, conformal prediction is computationally efficient and can be used with any prediction model.
Figure 1: Difference between exchangeability and spatial stationarity.
3.2 Proposed method: Local spatial conformal prediction (LSCP)
In spatial settings, data often exhibit significant dependence across locations, and taking the spatial dependence into account can improve the accuracy of prediction intervals. To account for this, it is advantageous to base predictions on nearby data points, as spatially proximate observations are likely to share similar distributions and therefore provide more reliable information. The recent study by (Barber et al., 2023) highlights the importance of weighting calibration data differently, depending on their relevance to the target prediction point.
To construct the prediction interval, we first split the dataset into a training set and a calibration set. The training set is used to train the prediction model $\hat{f}$, while the calibration set decides the width of the interval. We assume the calibration set consists of observations $(X(s_{1}),Y(s_{1})),\dots,(X(s_{n}),Y(s_{n}))$. For a new observation $(X(s_{n+1}),Y(s_{n+1}))$, the aim is to construct a prediction interval by selecting a neighborhood of data from the calibration data. Here we use $N(s_{n+1})$ to represent all the indices $i$ so that $s_{i}$ is in the neighborhood of $s_{n+1}$. The neighborhood can be determined via various criteria, and a common approach is to include all nearby points located within a specified distance threshold. In the paper, we use $k$-nearest neighbor for simplicity.
Given the trained model $\hat{f}$, the non-conformity scores are defined as $\hat{\varepsilon}(s_{i})=Y(s_{i})-\hat{f}(X(s_{i}))$. We then define the conditional cumulative distribution function (CDF) for the non-conformity scores based on the selected neighbors, denoted by
$F(e|\{\hat{\varepsilon}(s_{i})\}_{i\in N(s_{n+1})})=\mathbb{P}(\hat{\varepsilon}(s_{n+1})\leq e|\{\hat{\varepsilon}(s_{i})\}_{i\in N(s_{n+1})}).$
The conditional quantile $Q_{n}(p)$ is defined as
$Q_{n}(p)=\inf\{e\in\mathbb{R}:F(e|\{\hat{\varepsilon}(s_{i})\}_{i\in N(s_{n+1})})\geq p\}.$
The empirical quantile places an equal weight on all the points, which may not fully capture the dependence structure. Instead, we apply a quantile regression estimator $\hat{Q}_{n}$ on the residuals $\{\hat{\varepsilon}(s_{i})\}_{i\in N(s_{n+1})}$. The estimator, $\hat{Q}_{n}(\alpha)$, predicts the $\alpha$-quantile of the residual $\hat{\varepsilon}(s_{n+1})$ given the values of its neighbors. For computational efficiency, we use Quantile Random Forests from (Meinshausen & Ridgeway, 2006), although other quantile regression techniques could also be applied. The resulting prediction interval is:
$\hat{C}_{n}(X(s_{n+1}))=\left[\hat{f}(X(s_{n+1}))+\hat{Q}_{n}(\beta^{*}),\;\hat{f}(X(s_{n+1}))+\hat{Q}_{n}(1-\alpha+\beta^{*})\right],$ (1)
where $\beta^{*}=\operatorname{argmin}_{\beta\in[0,\alpha]}(\hat{Q}_{n}(1-\alpha+\beta)-\hat{Q}_{n}(\beta))$. Here $\beta$ is optimized to find the tightest interval.
To formalize this approach, let $\tilde{X}(s_{i})=(\hat{\varepsilon}(s_{j}))_{j\in N(s_{i})}$ and $\tilde{Y}(s_{i})=\hat{\varepsilon}(s_{i})$. The quantile regression, which learns the conditional quantile of $F(\tilde{Y}(s_{n+1})|\tilde{X}(s_{n+1}))$, computes:
$\hat{Q}_{n}(p)=\inf\{e\in\mathbb{R}:\sum_{i\in N(s_{n+1})}\omega_{i}\mathbf{1}\{\tilde{Y}(s_{i})\leq e\}\leq p\},$
where the weights $\omega_{i}$ are learned through quantile regression.
Figure 2: Illustration of Localized Spatial Conformal Prediction. Left: $k$-nearest neighbors with weights are used to predict the test point while providing uncertainty quantification. Right: The $k$-nearest neighborhood adapts to both dense and sparse observation regions. Red points represent test locations, with dashed circles marking neighborhoods of the 5-nearest neighbors. Darker grey shades indicate regions of lower uncertainty. LSCP leverages these neighboring points to construct prediction intervals at the test locations.
3.3 Comparison with related methods
3.3.1 Global Spatial Conformal Prediction (GSCP)
GSCP, introduced in (Mao et al., 2024), applies equal weighting to all non-conformity scores across the calibration dataset, where the non-conformity score is defined as
$\hat{\varepsilon}(s_{i})=\left|\frac{Y_{i}-\hat{f}(s_{i},X(s_{i}))}{\hat{\sigma}(s_{i},X(s_{i}))}\right|.$
The estimated quantile at any point is given by
$\hat{Q}_{n}(p)=\inf\{e\in\mathbb{R}:\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{\hat{\varepsilon}(s_{i})\leq e\}\leq p\}.$ (2)
The method can be viewed as an extension of split conformal prediction. Its primary advantage is the strong theoretical guarantee, achieved with minimal assumptions. When the spatial locations are sampled i.i.d. from a common distribution $F_{\varepsilon}(s)$, the data becomes exchangeable because switching the order does not change the joint probability. The marginal coverage automatically holds in this situation.
However, in real-world applications, the performance of global conformal methods like GSCP may be limited because data distributions can vary across locations. Unfortunately, the exchangeability does not hold for localized methods. Since GSCP lacks adaptivity to different locations, it tends to construct intervals that may be overly conservative to meet the marginal coverage requirement, leading to overcoverage in some regions.
Algorithm 1 Spatial Conformal Prediction
1:Dataset $\left\{\left(x(s_{i}),y(s_{i})\right)\right\}_{i=1}^{n}$, prediction algorithm $\mathcal{A}$, significance level $\alpha$
2:Prediction intervals $\{\widehat{C}_{n}\left(x(s_{n+1})\right)\}$
3:Split the dataset into training data and calibration data.
4:Train the prediction model $\hat{f}$ with the training data using the prediction algorithm $\mathcal{A}$.
5:Select the neighborhood of $s_{n+1}$ in the calibration data, denoted as $N(s_{n+1})$.
6:Compute the non-conformity scores $\hat{\varepsilon}(s_{i})=Y(s_{i})-\hat{f}(X(s_{i}))$ for all data in the calibration set.
7:Set $\tilde{Y}(s_{i})=\hat{\varepsilon}(s_{i})$ and $\tilde{X}(s_{i})=(\hat{\varepsilon}(s_{j_{1}}),\dots,\hat{\varepsilon}(s_{j_{|N(s_{i})|}}))$, where $s_{j}\in N(s_{i})$.
8:Fit quantile regression $\hat{Q}_{n}$ with all pairs $(\tilde{X}(s_{i}),\tilde{Y}(s_{i}))$ in the calibration data.
9:Obtain the prediction interval $\hat{C}_{n}(X(s_{n+1}))$.
3.3.2 Smoothed Local Spatial Conformal Prediction (SLSCP)
SLSCP, introduced in (Mao et al., 2024), improves upon previous GSCP by utilizing only nearby data to construct prediction intervals. The non-conformity scores are defined as in Equation 2, but the quantile is estimated locally:
$\hat{Q}_{n}(p)=\inf\{e\in\mathbb{R}:\sum_{i\in N(s_{n+1})}\omega_{i}\mathbf{1}\{\tilde{Y}(s_{i})\leq e\}\leq p\}.$ (3)
A key difference between our method and SLSCP is in the choice of weights $\omega_{i}$. In SLSCP, $\omega_{i}\propto k(\|s_{i}-s_{n+1}\|)$, where $k$ is a kernel function that depends solely on the distance between spatial locations. In contrast, our method learns $\omega_{i}$ through quantile regression based on features $\tilde{X}(s)$. This feature vector can contain additional information beyond spatial distance, allowing the learned weights to be more expressive and adaptive than a given kernel function. The numerical results in Section 5 demonstrate that our method significantly outperforms SLSCP in experiments.
Another key distinction lies in the theoretical foundations. Conditional on the location $s_{n+1}$, the asymptotic coverage of SLSCP depends on an infill sampling assumption, where data points become infinitely dense around $s_{n+1}$, requiring infinitely close data. Additionally, SLSCP assumes the data process to be a composition of an $L_{2}$-continuous spatial process and a locally i.i.d. noise process $\varepsilon(s)$. Under these assumptions, the process can be shown to be locally asymptotically exchangeable, achieving asymptotic coverage as data points accumulate around the test location.
In contrast, our method establishes a finite-sample bound for the conditional coverage gap of LSCP, which guarantees asymptotic coverage. A key distinction is that our result conditions on the feature $X(s_{n+1})$, rather than solely on the location $s_{n+1}$, making it more general and broadly applicable. We assume an additive data model with a stationary, spatially mixing error process rather than an i.i.d. process. Spatial mixing implies that dependence between data points diminishes with increasing distance. Instead of requiring infinitely close neighbors, we assume that the average dependence decreases as more neighbors are included. The assumption is reasonable in that, given a fixed dataset, selecting more neighbors for prediction increases the neighborhood size, leading to a natural decay in correlation as more distant neighbors are included. Furthermore, our setting can be extended to spatio-temporal settings, which are commonly encountered in real-world applications. In contrast, SLSCP cannot generalize to time-series data or similar scenarios, as the time index cannot become infinitely dense, as required by the infill sampling assumption.
3.3.3 Localized Conformal Prediction (LCP)
LCP, introduced in (Guan, 2023), provides a general framework for localized conformal prediction rather than the spatial setting only. The method combines GSCP and SLSCP in quantile estimation:
$\hat{Q}_{n}(p)=\inf\{e\in\mathbb{R}:\sum_{i=1}^{n}\omega_{i}\mathbf{1}\{\tilde{Y}(s_{i})\leq e\}\leq p\}.$ (4)
Similar to GSCP, LCP uses all calibration data for prediction, but like SLSCP, it applies weights to each data point. The weights are defined as $\omega_{i}\propto k(X(s_{i}),X(s_{n+1}))$, where $k$ is a user-specified kernel function measuring similarity between features. The choice allows for more flexibility than the location-based weights in SLSCP, as feature-based weights can capture more detailed information. However, LCP still relies on user-specified weights, meaning the chosen kernel function might not fully capture the dependence structure in the data.
In terms of theoretical assumptions, LCP assumes that the data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ are i.i.d., which ensures finite-sample marginal coverage. The assumption is stronger than all the other methods, which hinders the generality of the result.
Table 1: Comparison of assumptions and algorithms of three related localized conformal prediction methods.
LSCP (Ours) SLSCP ((Mao et al., 2024)) LCP ((Guan, 2023))
Algorithm Localized weighted quantile Localized weighted quantile Global weighted quantile
Weights Learned by quantile regression
Specifed kernel function on
location $s$
Specifed kernel function on
feature $X(s)$
Assumption on data distribution
Stationary and spatial mixing
noise $\varepsilon(s)$
Locally i.i.d. noise $\varepsilon(s)$ Globally i.i.d. data $(X(s),Y(s))$
Assumption on data model Additive model
Continuous mapping from
$L_{2}$-continuous spatial process
and noise process
No assumption
4 Theoretical results
4.1 Setting
Suppose the data is denoted by $\{Z(s_{i})\}_{i=1}^{n}$, where $Z(s)=(X(s),Y(s))$, $s\in\mathbb{R}^{d}$ denotes the spatial location, $X(s)\in\mathbb{R}^{d}$ is the feature vector and $Y(s)\in\mathbb{R}$ represents the univariate response. We assume that $Y(s)$ is generated from a true model with unknown additive noise:
$Y(s)=f(X(s))+\varepsilon(s),$ (5)
where $f$ is an unknown function and $\varepsilon(s)$ represents the noise process, whose marginal distribution is not necessarily Gaussian. Given a pretrained prediction model $\hat{f}$, we can compute the non-conformity scores
$\hat{\varepsilon}(s)=Y(s)-\hat{f}(X(s)).$
The estimated conditional distribution function $\widehat{F}(\varepsilon|x)$ is defined as
$\widehat{F}(\varepsilon|x)=\sum_{i=1}^{n}w_{t}(x)\mathbf{1}(\hat{\varepsilon}({s_{i}})\leq\varepsilon).$
Besides, we define the weighted empirical CDF as
$\widetilde{F}(\varepsilon|x)=\sum_{i=1}^{n}w_{t}(x)\mathbf{1}(\varepsilon({s_{i}})\leq\varepsilon).$
4.2 Preliminary
Stochastic Design: Random observation location.
The key distinction between time series and spatial data lies in the indexing of observations. For time series data, observations are indexed by a fixed, ordered sequence of time points, denoted as $(X_{t},Y_{t})$. This inherent ordering imposes a natural temporal structure, preventing data points from being exchangeable. In contrast, spatial data are indexed by locations that can be irregularly distributed across space, with each data point in a random field represented as $(X(s),Y(s))$, where $s$ denotes a spatial location. In stochastic design, the spatial locations $s$ are considered random, following some underlying distribution. If we further assume that the locations of both calibration and test data are independently sampled from the same underlying distribution, then exchangeability holds naturally. In this case, marginal coverage is guaranteed for global conformal prediction that constructs prediction region with the whole calibration dataset, as established in (Mao et al., 2024). However, this does not apply in time series contexts due to the fixed temporal order, making it impossible to freely exchange data points. In a spatial setting, local conformal methods violate this exchangeability assumption because they apply different weights or restrictions based on proximity. As a result, while global conformal methods achieve exchangeability and marginal coverage, local conformal methods require additional conditions to maintain validity. Interestingly, we can consider time series as a special case of spatial data by restricting the spatial domain to a single dimension and treating time as a scalar spatial index. In this case, each time point can be thought of as a distinct “location” on a one-dimensional grid, with the locations ordered sequentially. This perspective highlights that time series analysis is a subset of spatial analysis, albeit with the added constraint of temporal ordering.
Spatial mixing.
Next, we define the strong mixing co-efficient for random field $Z(\cdot)$. Let $\mathcal{F}_{Z}(T)=\sigma\langle Z(s):s\in T\rangle$ be the $\sigma$-field generated by the variables $\{Z(s)$ : $s\in T\},T\subset\mathbb{R}^{d}$. For any two subsets $T_{1}$ and $T_{2}$ of $\mathbb{R}^{d}$, let $\tilde{\alpha}\left(T_{1},T_{2}\right)=$ $\sup\left\{|\mathbb{P}(A\cap B)-\mathbb{P}A\mathbb{P}B|:A\in\mathcal{F}_{Z}\left(T_{1}\right),\quad B\in\mathcal{F}_{Z}\left(T_{2}\right)\right\}$, and let $d\left(T_{1},T_{2}\right)=\inf\left\{|x-s|:x\in T_{1},\quad s\in T_{2}\right\}$. For $d=1$, we define the strong mixing co-efficient as
$\alpha(a;b)=\sup\{\tilde{\alpha}((x-b,x],[y,y+b)):-\infty<x+a<y<\infty\},\quad a>0,b>0\text{. }$
Thus, $\alpha(a;\infty)$ corresponds to the standard strong mixing co-efficient commonly used in the time series case. To define the strong mixing co-efficient for $d\geq 2$, let $\mathcal{R}_{k}(b)\equiv\left\{\cup_{i=1}^{k}D_{i}:\sum_{i=1}^{k}\left|D_{i}\right|\leq b\right\}$ be the collection of all disjoint unions of $k$ cubes $D_{1},\ldots,D_{k}$ in $\mathbb{R}^{d},k\geq 1,b>0$. Following the definition in (Lahiri, 2003), the strong-mixing coefficient for the r.f. $Z(\cdot)$ for $d\geq 2$ is defined as
$\alpha(a;b)=\sup\left\{\tilde{\alpha}\left(T_{1},T_{2}\right):d\left(T_{1},T_{2}\right)\geq a,T_{1},T_{2}\in\mathcal{R}_{3}(b)\right\}.$
For simplifying the exposition, we further assume that there exist a nonincreasing function $\alpha_{1}(\cdot)$ with $\lim_{a\rightarrow\infty}\alpha_{1}(a)=0$ and a nondecreasing function $g(\cdot)$ such that the strong-mixing coefficients $\alpha(a,b)$ satisfies the inequality
$\alpha(a,b)\leq\alpha_{1}(a)g(b),\quad a>0,b>0,$ (6)
where the function $g(\cdot)$ is bounded for $d=1$ but may be unbounded for $d\geq 2$. Without loss of generality, we may assume that $\alpha_{1}(\cdot)$ is left continuous and $g(\cdot)$ is right continuous (otherwise, replace $\alpha_{1}(a)$ by $\alpha_{1}(a-)\geq\alpha_{1}(a)$ and $g(b)$ by $g(b+)\geq g(b))$. We shall specify exact conditions on the rate of decay of $\alpha_{1}(\cdot)$ and the growth rate of $g(\cdot)$ in the statements of the results below.
4.3 Assumptions
With the data $(X(s_{1}),Y(s_{1})),\cdots,(X(s_{n}),Y(s_{n}))$, we would like to construct a prediction region for $Y(s_{n+1})$ where only the feature $X(s_{n+1})$ is known. Define $\omega_{n}(x)$ as the weight function and $\widetilde{\omega}_{n}(X(s_{i}))$ to be the normalized weight over samples $s_{1},\cdots,s_{n}$ so that $\sum_{i=1}^{n}\widetilde{\omega}_{n}(X(s_{i}))=1$. The weight function can be some function that measures the similarity between $X(s)$ and $X(s_{n+1})$.
Assumption 4.1 (Weight decay).
There exists $\gamma>0$ so that the normalized weights satisfy
$\widetilde{\omega}_{n}(X(s_{i}))=o(n^{-\frac{1+\gamma}{2}}),$ (7)
for all $i$. In other word, $M_{n}=\max_{1\leq i\leq n}\widetilde{\omega}_{n}(X(s_{i}))=o(n^{-\frac{1+\gamma}{2}})$.
The requirement assumes that the normalized weights decay at a rate faster than $n^{-\frac{1}{2}}$. As we can see, $\widetilde{\omega}_{n}=\frac{1}{n}$ is a special case that satisfies this condition. The condition is also weaker than the requirement of $\widetilde{\omega}_{n}=O(\frac{1}{n})$ in a related study (Xu et al., 2024). Besides, the assumption can also be inferred from that of SLSCP (Mao et al., 2024) where a GBF kernel with infinitely close data leads to uniform weights.
Assumption 4.2 (Estimation quality).
There exists a sequence $\{\delta_{n}\}_{n\geq 1}$ such that
$\sum_{i=1}^{n}\|\hat{\varepsilon}(s_{i})-\varepsilon(s_{i})\|^{2}\leq\frac{\delta_{n}^{2}}{M_{n}},~{}~{}\|\hat{\varepsilon}(s_{n+1})-\varepsilon(s_{n+1})\|\leq\delta_{n}.$ (8)
The assumption requires that the average prediction error be bounded by a term $\delta_{n}^{2}$, which is a weaker requirement than the similar condition for bounding the average prediction error in (Xu & Xie, 2021). The reason lies in that $M_{n}$ is allowed to decay at a slower rate than $\frac{1}{n}$. Notably, our result of the coverage gap does not require $\delta_{n}$ to converge to zero. However, there are many cases where $\delta_{n}$ does indeed approach zero. For instance, extensive research has investigated the prediction error of neural networks. Under certain regularization conditions on $f$, (Barron, 1994) shows that $\delta_{n}=O\left(\frac{1}{\sqrt{n}}\right)$.
Assumption 4.3 (Stationary and spatial mixing).
The random field $\varepsilon(s)$ is stationary and strongly mixing with coefficient $\alpha$. The strong mixing coefficient can be bounded by $\alpha(a,b)\leq\alpha_{1}(a)g(b)$, where $\alpha_{1}$ is a nonincreasing function with $\lim_{a\rightarrow\infty}\alpha_{1}(a)=0$. We assume $E_{d\sim g_{n}}\alpha_{1}(d)^{2}\leq\frac{M}{n^{2}}$ where $g_{n}(d)$ is distribution of the distance between two sample points $s_{i}$ and $s_{j}$ ($1\leq i,j\leq n$). Besides, $F_{\varepsilon}(x)$ (the CDF of the true non-conformity score) is assumed to be Lipschitz continuous with constant $L_{n+1}>0$.
The requirement assumes that the true error $\varepsilon(s)$ is spatially stationary and strongly mixing, a condition weaker than the i.i.d. assumption used in (Mao et al., 2024) for spatial conformal prediction. Under this assumption, the original random field $\{X(s),Y(s)\}$ can still exhibit complex dependencies and be highly non-stationary. Here $b$ is a constant in the definition of spatial mixing, which can be any specified positive integer. The assumption further requires that the expectation of the mixing coefficient $\alpha_{1}(d)$ decays at a certain rate. The decay is reasonable in that a bigger $n$ implies sampling from a larger area, thereby increasing the average distance between calibration points. The assumption is analogous to a common condition in time series analysis, which requires the sum of strong mixing coefficients to be bounded. To sum up, this assumption indicates that dependence between data points diminishes as distance increases.
4.4 Coverage guarantee
With the assumptions, we can show that the distance between the empirical CDF of the residuals $\hat{\varepsilon(s)}$ and the noise $\varepsilon(s)$ can actually be bounded through Lemma 4.4
Lemma 4.4.
(Distance between the empirical CDF of $\varepsilon$ and $\hat{\varepsilon}$) Under Assumption 4.1 and 4.2,
$\sup_{x}|\widehat{F}_{n+1}(y)-\widetilde{F}_{n+1}(y)|\leq(L_{n+1}+1)\delta_{n}+2\sup_{y}|\widetilde{F}_{n+1}(y)-F_{\varepsilon}(y)|.$ (9)
Besides, we can also prove that the empirical CDF of the noise $\varepsilon(s)$ converges to its CDF with a high probability.
Lemma 4.5.
(Convergence of empicial CDF of $\varepsilon$) Under Assumptions 4.1-4.3, with probability higher than $1-(2+M+2\sqrt{M}g(b))n^{-\frac{2\gamma}{3}}(\log_{2}n+2)^{\frac{4}{3}}$,
$\sup_{y}\left|\widetilde{F}_{n+1}(y)-F_{\epsilon}(y)\right|\leq M_{n}n^{\frac{1+\gamma}{2}},$ (10)
where $\widetilde{F}_{n+1}(y)=\sum_{i=1}^{n}\widetilde{\omega}_{n}(X(s_{i}))\mathbf{1}(\epsilon(s_{i})\leq y)$.
Our main theorem is the following one, which establish the asymptotic convergence of conditional coverage using the previous lemma.
Theorem 4.6.
(Conditional Coverage) Under Assumption 4.1-4.3, for any $\alpha\in(0,1)$ and sample size $T$, we have
$\left|\mathbb{P}\left(Y(s_{n+1})\in\widehat{C}_{t-1}\left(X(s_{n+1})\right)\mid X(s_{n+1})\right)-(1-\alpha)\right|$ (11)
$\leq 4L_{n+1}\delta_{n}+6M_{n}n^{\frac{1+\gamma}{2}}+(4+2M+4\sqrt{M}g(b))n^{-\frac{2\gamma}{3}}(\log_{2}n+2)^{\frac{4}{3}}.$
We can establish the same result for marginal coverage with the tower law property.
Corollary 4.7.
(Marginal Coverage) Under Assumption 4.1-4.3, for any $\alpha\in(0,1)$ and sample size $T$, we have
$\left|\mathbb{P}\left(Y(s_{n+1})\in\widehat{C}_{t-1}\left(X(s_{n+1})\right)\right)-(1-\alpha)\right|$ (12)
$\leq 4L_{n+1}\delta_{n}+6M_{n}n^{\frac{1+\gamma}{2}}+(4+2M+4\sqrt{M}g(b))n^{-\frac{2\gamma}{3}}(\log_{2}n+2)^{\frac{4}{3}}.$
From Inequality LABEL:main_thm, we can see that the order of the coverage bound is controlled by $M_{n}n^{\frac{1+\gamma}{2}}$ and $n^{-\frac{2\gamma}{3}}(\log_{2}n+2)^{\frac{4}{3}}$. The first term is equal to $n^{\frac{\gamma-1}{2}}$ in the special case of $M_{n}=\frac{1}{n}$ and the second term diminishes when $n$ is large enough because $n$ is of higher order than $\log_{2}n$. As long as the estimation gap $\delta_{n}$ goes to zero when $n$ gets larger, the asymptotic conditional coverage can be inferred from the main theorem.
5 Experiments
In this section, we compare our proposed LSCP method against four baseline approaches. The first method, EnbPI ((Xu & Xie, 2021)), is designed for time series data and applies equal weight to the most recent observations. The second method, GSCP ((Mao et al., 2024)), uses the entire dataset for prediction, applying equal weights to all data points. The third method, SLSCP ((Mao et al., 2024)), leverages k-nearest neighbors for uncertainty quantification, assigning different weights to each point based on the spatial distance between the test data and the calibration data. The fourth method, LCP ((Guan, 2023)), utilizes all data points for prediction, weighting them according to the similarity between the features of test data and calibration data. Both SLSCP and LCP use the Gaussian kernel as a similarity measure.
We randomly split the dataset into three subsets: $40\%$ for training, $40\%$ for calibration, and $20\%$ for testing. For each method and dataset, the number of neighbors and the Gaussian kernel bandwidth are selected through $5$-fold cross-validation.
5.1 Synthetic data experiments
We begin by comparing our method, LSCP, with baseline methods across several simulated scenarios. For simplicity, we assume that the locations $s$ are uniformly sampled from the unit grid $[0,1]\times[0,1]$. The mean-zero stationary Gaussian process $X(s)$ is defined with a Matérn covariance function, having variance $\sigma^{2}=1$, range $\phi=0.1$, and smoothness $\kappa=0.7$. The scenarios are as follows:
  1. 1.
    $Y(s)=X(s)+\epsilon(s)$.
  2. 2.
    $Y(s)=X(s)|\epsilon(s)|$.
  3. 3.
    $Y(s)=X(s)+\sin(\|s\|_{2})\epsilon(s)$.
These scenarios include nonlinear and non-stationary settings that extend beyond the assumptions made in our theoretical framework. However, the empirical results demonstrate that our proposed LSCP method consistently outperforms other methods across all scenarios.
In Table 2, LSCP achieves the target coverage of $90\%$ across all cases while also maintaining a significantly narrower prediction interval. Figure 3 displays the interval width over the spatial domain for each scenario and the true residuals over the space. GSCP is omitted from these plots as it produces a uniform interval width across the space. A larger interval width indicates higher model uncertainty. These figures reveal that LSCP adapts more locally, while the other methods display smoother, more uniform patterns. Both EnbPI and SLSCP rely on k-nearest neighbors for constructing prediction intervals, differing primarily in their choice of weights. As bandwidth increases, the performance of SLSCP approximates that of EnbPI, since SLSCP weights approaching uniformity. A similar pattern is observed between LCP and GSCP.
To further assess the methods, we divide the unit grid into $10\times 10$ smaller areas, computing the coverage and average interval width for each method within each area. The resulting distributions are shown in Figure 5. LSCP and GSCP maintain an average coverage above the target $90\%$, with LSCP showing the most concentrated distribution, indicating stability and consistency across areas. In the average width plot, LSCP shows the smallest width while being the most consistent. On the other hand, LCP displays a similar performance compared to GSCP, and a major reason is that they both use the whole calibration dataset for prediction. When the bandwidth of the kernel grows larger, the weights in LCP converge to $\frac{1}{n}$ in GSCP. Overall, our proposed LSCP not only provides tighter valid regions but also demonstrates the highest consistency.
Scenario 1
Scenario 2
Scenario 3
(a) Residual
(b) LSCP
(c) EnbPI
(d) SLSCP
(e) LCP
Figure 3: The heatmaps illustrate the width of the prediction intervals for each method across the three scenarios.
Table 2: Simulation: The table presents a comparison of the coverage and prediction interval width for five methods across three scenarios. LSCP consistently achieves the target coverage of $90\%$ in all cases, while also maintaining the smallest prediction interval width. The number in the bracket shows the bandwidth selected by $5$-fold cross-validation. When the bandwidth is large, the performance of LCP is similar to GSCP, and the performance of SLSCP is close to EnbPI.
Method S1 Coverage S1 Width S2 Coverage S2 Width S3 Coverage S3 Width
LSCP 92% 1.44 90.4% 0.64 91.3% 0.97
EnbPI 88.5% 1.83 89.2% 0.75 86.2% 1.12
GSCP 89.4% 1.92 89.7% 0.77 89.6% 1.30
SLSCP 90.6% ($0.03$) 1.64 88.2% ($0.03$) 0.73 89.1% ($0.03$) 1.07
LCP 89.6% ($0.5$) 1.92 94.1% ($0.02$) 0.78 91.8% ($0.01$) 1.43
(a) New Mexico
(b) Georgia
Figure 4: Mobile signal score from Ookla dataset. The plots show the distribution of data accross the state.
5.2 Real data experiments
In the real data experiments, we utilize the mobile network measurement data from the Ookla public dataset, which records user-reported statewide mobile internet performance. Our analysis focuses on the New Mexico and Georgia dataset. Figure 4 illustrates the spatial distribution of data across the two states. As we can see, the data distributes unevenly across urban and suburb areas. We employ kernel regression as the prediction method $\hat{f}$, and compare the performance of the conformal methods. The New Mexico dataset contains $24983$ observations in total and the Georgia dataset has $28587$ data points.
As summarized in Table 3, our proposed LSCP method demonstrates a clear advantage over competing methods, offering both significantly narrower prediction intervals and superior coverage. Similar to the synthetic experiments, we divide each state into $10\times 10$ uniform grids and calculate the coverage and average interval width for the test data within each grid. These metrics allow us to examine the spatial consistency of the methods. The violin plots in Figure 5 visually represent the distribution of coverage and interval width across the grids. It is important to note that the averages presented in the violin plots differ from those in Table 3. While the table reports averages over all test points, the violin plots reflect the average of grid areas. Consequently, the violin plots offer a different perspective on how the methods perform spatially, highlighting whether they behave uniformly across the entire state. The coverages reported in the table are generally higher because these methods perform well in the urban areas, where data is more concentrated.
The results reveal that LSCP consistently achieves high coverage while maintaining a narrow and spatially uniform interval width, outperforming alternative approaches. This is particularly evident in the violin plots, where LSCP exhibits one of the highest and most consistent coverage rates and the most uniform interval widths across the state. This underscores the robustness and adaptability of LSCP to non-uniform data distributions, offering reliable and geographically consistent prediction intervals. These findings align with those from our synthetic study, further validating the effectiveness of LSCP in diverse and challenging data settings.
Scenario 1
Scenario 2
Scenario 3
New Mexico
Georgia
(a) Coverage
(b) Width
Figure 5: The violin plots on the left show the distribution of coverage across different areas, while the plots on the right show the distribution of width. Each row represents a different scenario.
Table 3: Real data: The table presents a comparison of the coverage and prediction interval width for five methods on mobile signal data. LSCP reaches target coverage with the smallest prediction interval width. Since the bandwidth chosen by CV is large, the performance of LCP is similar to GSCP, and the performance of SLSCP is close to EnbPI.
Method NM Coverage NM Width GA coverage GA width
LSCP 92.6% 211.3 90.6% 130.8
EnbPI 88.4% 272.2 88.6% 166.3
GSCP 90.1% 295.9 89.2% 167.3
SLSCP (bd=0.2) 89.8% 276.8 89.6% 167.8
LCP (bd=0.1) 89.7% 266.4 89.3% 166.7
6 Discussion
This paper introduces the Localized Spatial Conformal Prediction (LSCP) method, which addresses key limitations in spatial and spatio-temporal prediction for non-exchangeable data. Traditional methods, such as Kriging, often depend on strong parametric assumptions that may not hold in complex real-world scenarios, including heterogeneous or large-scale spatial datasets. While machine learning techniques offer flexibility and predictive power, they typically lack robust uncertainty quantification, which is critical for applications requiring reliable decision-making.
Our results show that LSCP significantly outperforms existing methods, including GSCP, SLSCP, and EnbPI, by achieving more accurate coverage and narrower prediction intervals. This advantage is particularly pronounced in scenarios with non-stationary and non-Gaussian data, where LSCP’s flexibility enables it to effectively handle such complexities. Moreover, LSCP’s theoretical framework supports extension to spatio-temporal settings where exchangeability is unlikely to hold.
In synthetic experiments, LSCP consistently meets the target coverage with narrower intervals, even in cases beyond its assumptions, demonstrating robustness across diverse scenarios. Real-world experiments further validate LSCP’s utility, showing that it generates detailed uncertainty maps with tighter intervals than baseline methods. These features underscore LSCP’s potential as a reliable and scalable tool for spatial uncertainty quantification, offering both theoretical guarantees and practical performance benefits.